OpenStudy (anonymous):

use natural log to solve 3^2x=20 (its a multiple choice problem) the possible answers are a. ln20/ln6 b. ln(10/3) c. ln20/ln9 or d. ln20-ln3/2

OpenStudy (kirbykirby):

\(3^{2x}=20\) log both sides: \(\ln(3^{2x})=\ln 20\) \(2x \ln 3 = \ln 20\) \[2x = \frac{\ln 20}{\ln 3}\] \[ x=\frac{\ln 20}{2 \ln 3}=\frac{\ln 20}{\ln3^2}=\frac{\ln 20}{\ln 9}\]

OpenStudy (anonymous):

Thank you so much!!! I get it now :) you are a lifesaver!

OpenStudy (kirbykirby):

:)